Does there exist a plane targent to
x^2 −2y^2 +2z^2 = 8 and which passes through 2x+3y+2z = 8, x−y+2z = 5? Justify your answer.
1
Expert's answer
2020-05-18T18:49:14-0400

Given "f(x,y,z)=x^{2}-2y^{2}+2z^{2}-8=0"  with point "P(a,b,c)" .

"\implies a^{2}-2b^{2}+2c^{2}-8=0"

Now "(f_{x},f_{y},f_{z})=(2x,-4y,4z)" and so at point P it is: "(2a,-4b,4c)" .

So, Equation of the tangent plane is "2a(x-a)-4b(y-b)+4c(z-c)=0"

"\implies ax−2by+2cz=8"

The system "2x+3y+2z = 8, x-y+2z=5"  is a line, we choose points on the line.

The points (3,0,1) and "(-1,1,\frac{7}{2})"  of the given line must lie in this tangent plane,

then we have "c=\frac{8-3a}{2}"​ and "b=\frac{40-23a}{4}" .

Using the equation "a^{2}-2b^{2}+2c^{2}-8=0"

"\implies a^2-\frac{\left(40-23a\right)^2}{8}+\frac{\left(8-3a\right)^2}{2}-8\:=0",

"\implies -485a^2+1648a-1408=0"

"\implies a=\frac{824}{485}-i\frac{8\sqrt{61}}{485},\:a=\frac{824}{485}+i\frac{8\sqrt{61}}{485}"

we do not find a real solution for a.

Answer: The tangent plane does not exist.